# On Minimal Achievable Quotas in Multiwinner Voting

### v1

@@ Abstract
Justified representation (JR) and extended justified representation (EJR) are well-established proportionality axioms in approval-based multiwinner voting. Both axioms are always satisfiable, but they rely on a fixed quota (typically Hare or Droop), with the Droop quota being the smallest one that guarantees existence across all instances. With this observation in mind, we take a first step beyond the fixed-quota paradigm and introduce proportionality notions where the quota is instance-dependent. We demonstrate that all commonly studied voting rules can have an additive distance to the optimum of $\frac{k^{2}}{(k+1)^{2}}$. Moreover, we look into the computational aspects of our instance-dependent quota and prove that determining the optimal value of $\alpha$ for a given approval profile satisfying $\alpha$-JR is NP-complete. To address this, we introduce an integer linear programming (ILP) formulation for computing committees that satisfy $\alpha$-JR, and we provide positive results in the voter interval (VI) and candidate interval (CI) domains.
@@ 1 Introduction
In the context of multiwinner approval voting, fairness is typically understood in terms of proportionality: the elected committee should reflect, to an appropriate extent, the fraction of voters who share similar preferences or approval sets. This guarantees that sufficiently large groups of voters receive adequate representation, while preventing smaller groups from being disregarded. Proportionality axioms, like justified representation (JR) and extended justified representation (EJR), formalize this intuition and have become a central concept in the study of fairness in multiwinner elections. Given the size of the voter set $n$ and committee size $k$, a group of voters $S$ of size at least $n/k$ that agrees on a common candidate is called cohesive and should be able to demand some form of representation in the committee. The Hare quota (=n/k=n/k) is deeply ingrained in the definitions of most proportionality axioms in multiwinner voting — including JR and EJR, but also proportional justified representation (PJR) and full justified representation (FJR).
Nevertheless, there is a variety of situations where these proportionality axioms fail to capture intuitive fairness. JR demands that every cohesive group has at least one voter $v\in S$ who approves a candidate in the committee. Consider, for example, the election instance provided in Figure˜1. To be cohesive, the group has to contain at least $5$ voters. It should be easy to see that the committee $W=\{c_{3},c_{4}\}$ satisfies JR, and, thereby, disregards the groups $\{1,2,3,4\}$ and $\{7,8,9,10\}$. These groups fall just short of meeting the quota. This instance, however, permits a more balanced representation of all voters by selecting the committee $W^{\prime}=\{c_{1},c_{2}\}$.
@@
Figure 1: A voting instance with $N=[10]$, $C=\{c_{1},c_{2},c_{3},c_{4}\}$ and $k=2$ in which any committee containing $c_{3}$ satisfies JR. Voters are denoted by an integer and precisely approve the candidates stacked above them.
This phenomenon extends to all fairness axioms that are defined with respect to these static quota requirements. Even the core is not immune to this issue. [^1] For a formal definition, consider the paper by Peters et al. (2021b). Figure˜2 illustrates an example in which the committee $\{c_{1},c_{2},c_{3}\}$ belongs to the core, even though voters 551212 remain unrepresented. The three voters supporting $c_{4}$ once again fall short of the quota and can therefore not form a valid deviating coalition.
@@
Figure 2: A voting instance with $N=[12]$, $C=\{c_{1},\dots,c_{7}\}$ and $k=3$ in which the committee $W=\{c_{1},c_{2},c_{3}\}$ lies within the core even though a large group of voters remains unrepresented (example taken from Peters et al. (2021b)).
@@
"[P]roperties like the core (and also EJR, PJR, and JR) prevent specific pathological situations, but beyond their definitions, do not provide intuitive justifications for why a committee they allow should be selected" — Peters et al. (2021b).
For some proportionality axioms, representation cannot be guaranteed for groups smaller than the Droop quota ($\frac{n}{k+1}$) (Droop, 1881). Hence, fixed quotas may yield examples like that in Figure˜1 or render the axiom unsatisfiable. This raises the question of whether, instead of relying on a uniform quota across all instances, we should make it adaptive with respect to the preference profile. Our experiments show that the quota values for which JR or EJR remains satisfiable tend to be much smaller than $\frac{n}{k+1}$ A similar observation holds for the ordinal case (Bardal et al., 2025).
In this paper, we propose variable quotas. The central idea is to reduce the quota $n/k$ by a factor $\alpha$, approaching the infimum value at which JR or EJR ceases to be feasible. We refer to this approach as the $\alpha$-extension of cohesiveness. For the case of JR, finding the optimal $\alpha$-value corresponds to minimizing the largest group of unrepresented voters approving a common candidate.
To illustrate this concept more concretely, consider the following example: A company has four locations, each comprising five members. The task is to form a representative committee of three members. Within each location, all members approve a location representative aware of the problems at this location. Each location has a manager, salesperson, and engineer, with each role group approving a role representative. The remaining two members per location are working students, who only approve their location representative. Known approval-based multiwinner rules like PAV or CC elect three of the four location representatives as they receive the broadest approval support. However, this outcome fails to represent an entire location. We argue that a more proportionally fair outcome — and the optimal committee with respect to $\alpha$-JR — would instead include the three group representatives, ensuring that no large and homogeneous segment of the electorate, like an entire location of a company, remains unrepresented.
@@
In this paper, we introduce a class of proportionality axioms, denoted $\alpha$-JR and $\alpha$-EJR, which generalize the existing proportionality axioms JR and EJR. There are instances for which existing voting rules perform very badly with respect to these axioms. More specifically, we demonstrate that all commonly studied voting rules can produce committees with an additive distance of $\frac{k^{2}}{(k+1)^{2}}$ away from the optimal $\alpha$-value. Next to that, we show that computing the optimal $\alpha$-value for a voting instance is NP-hard, even for JR. In order to find the optimal value for JR, we provide an ILP. When considering restricted voting instances, we can show positive results: In the voter- and candidate-interval domain, we can efficiently compute the optimal committee with respect to $\alpha$-JR. For party-list profiles, we provide an algorithm to construct an optimal $\alpha$-EJR committee in polynomial time. Lastly, we provide empirical evaluation on the distribution of the optimal $\alpha$-value for synthetically generated instances and show that existing voting rules can, to some extent, approximate the optimal $\alpha$-value on average.
@@ 2 Related Work
Our work builds on the growing literature on fairness and proportionality in approval-based multiwinner voting (see, e.g., (Lackner and Skowron, 2023; Aziz et al., 2017; Sánchez-Fernández et al., 2017; Skowron et al., 2017a)). The central idea in this line of research is that voters who agree on a set of candidates should receive influence in the elected committee proportional to their size. The notions of JR and EJR, introduced by Aziz et al. (2017), translate this into the setting of multiwinner approval voting and, thereby, mark the starting point of proportionality in approval-based multiwinner voting. Building on these axioms, subsequent work has extended proportionality axioms and analyzed their implications for prominent rules (Brill and Peters, 2023; Brill et al., 2024). These ideas have also been generalized beyond multiwinner elections to participatory budgeting (Peters et al., 2021a), and general voting models with feasibility constraints (Masařík et al., 2024).
One aspect that has received relatively little attention is the role of quotas, i.e., the requirement that a group of like-minded voters must reach a certain threshold to secure representation in the committee. The most common quota that has been used in the literature so far is the Hare quota ($\frac{n}{k}$). For this, it has been shown that JR and EJR are always satisfiable (Aziz et al., 2017). This also holds for some stronger versions of the axioms, like EJR+ and FJR. More recently, these representation axioms have been considered for the Droop quota ($\frac{n}{k+1}$) (Casey and Elkind, 2025). In this paper, the authors demonstrate that proportionality axioms based on the Droop quota can be satisfied for any given voting instance. They show that many voting rules and concepts originally developed for the Hare quota can be effectively adapted to the Droop quota setting. Before this work, the Droop quota was primarily studied within the apportionment framework (Pukelsheim, 2017; Brill et al., 2020).
@@
Closest related to ours is the extensive work by Janson (2018), who studied and surveyed thresholds of proportionality. More specifically, he asked the question of the smallest proportion of voters that can be guaranteed a good outcome, where a good outcome corresponds to committees satisfying various proportionality notions. While he mostly focused on fixing the proportion of votes achievable by sequential Phragmén and Thiele rules, we consider more voting rules and look into approaches and results around the computational aspects of minimizing the quota.
@@
More recently, Bardal et al. (2025) developed quantitative measures of proportionality in the ordinal setting, thereby strengthening the proportionality guarantees of so-called solid coalitions. Their work primarily focused on empirical analyses of real-world data, whereas our study centers around the theoretical aspects of optimizing the $\alpha$-value in the approval setting and exploring results in the voter- and candidate-interval domain. In parallel to us, they introduced the parameter $\alpha$, which for values less than $1$ makes proportionality for solid coalitions more demanding. Skowron et al. (2017b) formalized proportionality in the ranking setting when voters have approval preferences and introduced a measure of group satisfaction within this setting.
@@
Alternative approaches to capturing fairness in multiwinner voting have been explored beyond JR–based axioms (Brill and Peters, 2024; Kehne et al., 2025; Lackner and Skowron, 2020). Brill et al. (2025) introduced the axiom of individual representation (IR), which strengthens the earlier concept of semi-strong justified representation proposed by Aziz et al. (2017). While semi-strong JR requires that every sufficiently large and cohesive group of voters is represented in the committee at least once, IR refines this principle by assigning each voter a guarantee relative to the largest $\ell$-cohesive group she is able to form for a given profile. IR is not always satisfiable: there exist profiles for which no committee provides individual representation, and non-trivial approximation guarantees are unattainable without imposing restrictions on the domain of admissible profiles. Kehne et al. (2025) changed the viewpoint of fairness to be more candidate-centric. Their desideratum of fairness says that candidates with similar sets of supporters should receive similar representation.
@@
Lastly, we mention the proportionality degree (PD) as a different way of quantifying proportionality in multiwinner approval voting. Informally, this is a function that specifies the average satisfaction of a voter in an $\ell$-cohesive group. This approach strictly enforces the quota while allowing greater flexibility in determining the average satisfaction level of each group. Skowron (2021) used this to measure the proportionality guarantees that different voting rules can achieve. He showed that the sequential variant of the proportional approval voting rule gets higher PD than sequential Phragmén. Janeczko and Faliszewski (2022) looked into computational aspects of the proportionality degree in committee elections. They show that deciding if a committee with a given PD exists is NP-hard and that verifying if a given committee provides a given PD is coNP-complete.
@@ 3 Preliminaries
We consider the standard approval-based multiwinner voting framework. Let $N=\{v_{1},\dots,v_{n}\}$ be the set of voters, $C=\{c_{1},\dots,c_{m}\}$ be the set of candidates, and $k\in\operatorname{\mathbb{N}}$ be the committee size with $k\leq m$. For every voter $v\in N$, we denote the set of candidates she approves of as $A_{v}\subseteq C$, and we write $N_{c}=\{v\in N:c\in A_{v}\}$ for the set of supporters of a candidate $c\in C$. A voting instance $I$ is defined by an approval profile $A=(A_{v})_{v\in N}$ and the committee size $k$, i.e., $I=(A,k)$. Finally, a committee $W$ is a subset of the candidates with cardinality $k$. In order to determine which committee should be selected, different multiwinner voting rules have been proposed. In this paper, we consider commonly used approval-based multiwinner voting rules, like the method of equal shares (MES), sequential Phragmén (seq-Phragmén), Chamberlin-Courant (CC), proportional approval voting (PAV), and greedy justified candidate rule (GJCR). We refer the reader to Appendix˜A for the formal definitions of these rules.
@@ 4 Quota-Based Strengthening of Proportionality
As stated in the beginning, most proportionality axioms rely on the principle that a group is entitled to representation only if it satisfies some cohesiveness condition. Cohesiveness in its general form demands that a voter group satisfies the Hare quota (Aziz et al., 2017; Sánchez-Fernández et al., 2017).
@@
For $\ell\in\operatorname{\mathbb{N}}_{+}$, a group $S\subseteq N$ is $\ell$-cohesive if (1) $\lvert S\rvert\geq\ell\cdot\frac{n}{k}$ and (2) $\lvert\bigcap_{v\in S}A_{v}\rvert\geq\ell$.
In this work, we introduce a more general version of this definition, which introduces a parameter $\alpha$, making the quota requirement dynamic. The parameter $\alpha$ acts as a scaling factor that adjusts the group size required for representation. Smaller $\alpha$ values correspond to a more demanding standard of proportionality, as even smaller groups can claim representation.
@@
For $\ell\in\operatorname{\mathbb{N}}_{+}$ and $\alpha\in\operatorname{\mathbb{R}}_{\geq 0}$, a group $S\subseteq N$ is $(\alpha,\ell)$-cohesive if (1) $\lvert S\rvert\geq\alpha\cdot\ell\cdot\frac{n}{k}$ and (2) $\lvert\bigcap_{v\in S}A_{v}\rvert\geq\ell$.
Remark 1.
Changing the quota size has, up until now, mostly been done in the other direction, i.e., one relaxes the quota to make a proportionality axiom easier to satisfy (Dong et al., 2025a, b; Jiang et al., 2020).
With this generalized cohesiveness definition, let us now modify existing proportionality axioms:
@@
A committee $W$ satisfies $\alpha$-EJR+ if there does not exist a candidate $c\in C\setminus W$, a group of voters $S\subseteq N$, and $\ell\in\operatorname{\mathbb{N}}_{+}$ with $|S|\geq\alpha\cdot\ell\cdot\frac{n}{k}$ such that
@@
Clearly, for the case $\alpha=1$, the three proportionality axioms correspond to the original notions of JR, EJR, and EJR+. Therefore, several voting rules are known to provide guarantees of satisfiability. More recently, it has been established that for any given instance, there exists an $\alpha<1$ such that $\alpha$-EJR+ still can be satisfied.
@@
In this paper, we are interested in the minimal possible quota that is still achievable by a given instance. Therefore, we like to find the infimum $\alpha$-value that still allows for an instance to satisfy $\alpha$-$\Phi$ where $\Phi\in\{\text{JR},\text{EJR},\text{EJR+}\}$. That being said, let
@@
We refer to $\alpha_{\Phi}(W)$ as the $\alpha$-value of a committee $W$ with respect to the proportionality axiom $\Phi$. The minimum achievable $\alpha$-value for an instance $I=(A,k)$ is denoted by
$\alpha^{*}_{\Phi}(I)=\min_{W\subseteq C:\lvert W\rvert=k}\alpha_{\Phi}(W).$
We also refer to this as the optimal $\alpha$-value for an instance $I$ with respect to $\Phi$. If the instance $I$ is clear from the context, we omit it and just write the optimal $\alpha$-value with respect to $\Phi$.
The notion of $\alpha_{\Phi}$ is equivalent to finding the largest $\alpha$-value for which there is a committee $W$ violating $\alpha$-$\Phi$. Observe that by definition, the value is greater than or equal to 0. If all voters with a non-empty ballot approve a candidate in the committee $\alpha_{\text{JR}}=0$, and if $\lvert A_{v}\cap W\rvert\geq\min(\lvert A_{v}\rvert,k)$ for all $v\in N$, then $\alpha_{\text{EJR}}=0$.

### v2

@@ Abstract
Justified representation (JR) and extended justified representation (EJR) are well-established proportionality axioms in approval-based multiwinner voting. Both axioms are always satisfiable, but they rely on a fixed quota (typically Hare or Droop), with the Droop quota being the smallest one that guarantees existence across all instances. With this in mind, we take a step beyond the fixed-quota paradigm by studying instance-dependent proportionality notions. More specifically, we minimize the quota requirements for JR and EJR using the parameter $\alpha$. We demonstrate that all commonly studied voting rules can have an additive gap to the optimum of $\frac{k^{2}}{(k+1)^{2}}$. Moreover, we examine the computational aspects of our instance-dependent quota and prove that determining the optimal value of $\alpha$ for a given approval profile that allows some committee to satisfy $\alpha$-JR is NP-complete. To address this, we introduce an integer linear programming (ILP) formulation for computing committees that satisfy $\alpha$-JR, and we provide positive computational results in the voter interval (VI) and candidate interval (CI) domains.
@@ 1 Introduction
In the context of multiwinner approval voting, fairness is typically understood in terms of proportionality: the elected committee should reflect, to an appropriate extent, the fraction of voters who share similar approval sets. This guarantees that sufficiently large groups of voters receive adequate representation, while preventing smaller groups from being disregarded. Proportionality axioms, like justified representation (JR) and extended justified representation (EJR), formalize this intuition and have become a central concept in the study of fairness in multiwinner elections. Given the size of the voter set $n$ and committee size $k$, a group of voters $S$ of size at least $n/k$ that agrees on a common candidate is called cohesive and should be able to demand some form of representation in the committee. The Hare quota (=n/k=n/k) is deeply ingrained in the definitions of most proportionality axioms in multiwinner voting, including JR and EJR, as well as proportional justified representation (PJR) and full justified representation (FJR).
Nevertheless, there are various situations where these proportionality axioms fail to capture intuitive fairness. JR demands that every cohesive group $S$ has at least one voter $v\in S$ who approves a candidate in the committee. Consider, for example, the election instance provided in Figure˜1. For the group $S$ to be cohesive, it must contain at least $5$ voters. It should be easy to see that the committee $W=\{c_{3},c_{4}\}$ satisfies JR, and, thereby, disregards the groups $\{1,2,3,4\}$ and $\{7,8,9,10\}$. These groups fall just short of meeting the quota. This instance, however, permits a more balanced representation of all voters by selecting the committee $W^{\prime}=\{c_{1},c_{2}\}$.
@@
Figure 1: A voting instance with 10 voters, candidate set $\{c_{1},c_{2},c_{3},c_{4}\}$ and $k=2$ in which any committee containing $c_{3}$ satisfies JR. Voters are denoted by an integer and approve the candidates stacked above them.
This phenomenon extends to all fairness axioms defined in relation to these static quota requirements. Even the core is not immune to this issue. [^1] For a formal definition, consider the paper by Peters et al. (2021b). Figure˜2 illustrates an example in which the committee $\{c_{1},c_{2},c_{3}\}$ belongs to the core, even though voters 551212 remain unrepresented. The three voters supporting $c_{4}$ once again fall short of the quota and can therefore not form a valid deviating coalition.
@@
Figure 2: A voting instance with 12 voters, candidate set $\{c_{1},\dots,c_{7}\}$ and $k=3$ in which the committee $W=\{c_{1},c_{2},c_{3}\}$ lies within the core even though a large group of voters remains unrepresented (example taken from (Peters et al., 2021b)).
@@
"Properties like the core (and also EJR, PJR, and JR) prevent specific pathological situations, but beyond their definitions, do not provide intuitive justifications for why a committee they allow should be selected" — Peters et al. (2021b).
For some proportionality axioms, representation cannot be guaranteed for groups smaller than the Droop quota ($\frac{n}{k+1}$) (Droop, 1881). Hence, fixed quotas may yield examples like that in Figure˜1 or render the axiom unsatisfiable. This raises the question of whether, instead of relying on a uniform quota across all instances, one should adapt it with respect to the preference profile. Our experiments show that the quota values for which JR or EJR remain satisfiable tend to be much smaller than $\frac{n}{k+1}$. A similar observation has been made for the ordinal case (Bardal et al., 2025).
This paper analyzes variable quotas, contributing to and expanding upon the existing literature (Janson, 2018; Bardal et al., 2025). The central idea is to reduce the quota $n/k$ by a factor $\alpha$, approaching the infimum value at which JR or EJR ceases to be feasible. In the case of JR, finding the optimal $\alpha$-value corresponds to minimizing the size of the largest group of unrepresented voters who approve a common candidate.
To illustrate this concept more concretely, consider one more example: A company has four locations, each comprising five members. The task is to form a representative committee of three members. Within each location, all members approve a location representative who is aware of the problems at that location. Each location has a manager, a salesperson, and an engineer, with each role group approving a representative for their role. The remaining two members per location are working students, who only approve their location representative. Known approval-based multiwinner rules, such as PAV or CC, elect three of the four location representatives because they receive the broadest approval support. However, this outcome fails to represent an entire location. We argue that a more proportionally fair outcome — and the optimal committee with respect to $\alpha$-JR — would instead include the three group representatives, ensuring that no large and homogeneous segment of the electorate, such as an entire company location, remains unrepresented.
@@
In this paper, we study a class of proportionality axioms, denoted $\alpha$-JR and $\alpha$-EJR, which generalize the existing proportionality axioms JR and EJR. There are instances for which existing voting rules perform very badly with respect to these axioms. More specifically, we demonstrate that all commonly studied voting rules can produce committees with an additive gap to the optimal $\alpha$-value of $\frac{k^{2}}{(k+1)^{2}}$. Additionally, we show that computing the optimal $\alpha$-value for a voting instance is NP-hard, even for JR. To find the optimal value for JR, we formulate an ILP. Further, in the VI and CI domains, we can efficiently compute the optimal committee with respect to $\alpha$-JR. For party-list profiles, we provide an algorithm that constructs an optimal $\alpha$-EJR committee in polynomial time, corresponding to the d’Hondt method in the apportionment setting. Lastly, we provide empirical evaluation on the distribution of the optimal $\alpha$-value for synthetically generated instances and show that existing voting rules can, to some extent, approximate the optimal $\alpha$-value on average.
Setting $\alpha$-JR $\alpha$-EJR $\alpha$-EJR+
Computing $\alpha_{\Phi}^{*}(I)$ NP-complete (Thm 5.2) NP-hard (Thm 5.2) NP-complete (Thm 5.2)
Computing $\alpha_{\Phi}^{W}(I)$ $O(nm)$ (Thm 5.1) coNP-hard (Thm 5.4) $O(mnk)$ (Thm 5.5)
Restricted domains (computing $\alpha_{\Phi}^{*}(I)$)
[…]
VI domain $O(n^{2}m\log n)$ (Thm 6.5) ? ?
CI domain $O(m^{2}n\log n)$ (Thm 6.8) ? ?
Table 1: Overview of our computational results. Cells with a question mark indicate that the question is still open.
@@ 2 Related Work
Our work builds on the growing literature on fairness and proportionality in approval-based multiwinner voting (see, e.g., (Lackner and Skowron, 2023; Aziz et al., 2017; Sánchez-Fernández et al., 2017; Skowron et al., 2017a)). The central idea in this line of research is that voters who agree on a set of candidates should receive influence in the elected committee proportional to their size. The notions of JR and EJR, introduced by Aziz et al. (2017), translate this into the setting of multiwinner approval voting, thereby marking the starting point of proportionality in approval-based multiwinner voting. Building on these axioms, subsequent work has extended proportionality axioms and analyzed their implications for prominent rules (Brill and Peters, 2023; Brill et al., 2024). These ideas have also been generalized beyond multiwinner elections to participatory budgeting (Peters et al., 2021a), and general voting models with feasibility constraints (Masařík et al., 2024).
One aspect that has received relatively little attention is the role of quotas, i.e., the requirement that a group of like-minded voters must reach a certain threshold to secure representation in the committee. The most common quota used in the literature is the Hare quota ($\frac{n}{k}$). For this, it has been shown that JR and EJR are always satisfiable (Aziz et al., 2017). This also holds for some stronger versions of the axioms, like EJR+ and FJR. More recently, these representation axioms have been considered for the Droop quota ($\frac{n}{k+1}$) (Casey and Elkind, 2025). In this paper, the authors demonstrate that proportionality axioms based on the Droop quota can be satisfied for any given voting instance. They show that many voting rules and concepts originally developed for the Hare quota can be effectively adapted to the Droop quota setting. Before this work, the Droop quota was primarily studied within the framework of apportionment (Pukelsheim, 2017; Brill et al., 2020).
@@
The work closest to ours is that of Janson (2018), who studied and surveyed thresholds of proportionality. More specifically, he asked for the smallest proportion of voters that can be guaranteed a good outcome, where a good outcome corresponds to committees satisfying various proportionality notions. While he mostly focused on adjusting the proportion of votes achievable by sequential Phragmén and Thiele rules, we consider additional voting rules and examine computational approaches and results related to minimizing the quota.
@@
More recently, Bardal et al. (2025) developed quantitative measures of proportionality in the ordinal setting, thereby strengthening the proportionality guarantees of so-called solid coalitions. Their work primarily focused on empirical analyses of real-world data, whereas our study centers on the theoretical aspects of optimizing the $\alpha$-value in the approval setting and on exploring results in the voter- and candidate-interval domains. In parallel to us, they study the parameter $\alpha$, which for values less than $1$ makes proportionality for solid coalitions more demanding. Skowron et al. (2017b) formalized proportionality in the ranking setting when voters have approval preferences and introduced a measure of group satisfaction within this setting.
@@
Alternative approaches to capturing fairness in multiwinner voting have been explored beyond JR–based axioms (Brill and Peters, 2024; Kehne et al., 2025; Lackner and Skowron, 2020). Brill et al. (2025) introduced the axiom of individual representation (IR), which strengthens the earlier concept of semi-strong justified representation proposed by Aziz et al. (2017). While semi-strong JR requires that every sufficiently large and cohesive group of voters is represented in the committee at least once, IR refines this principle by assigning each voter a guarantee relative to the largest $\ell$-cohesive group she is able to form for a given profile. IR is not always satisfiable: there exist profiles for which no committee provides individual representation, and non-trivial approximation guarantees are unattainable without imposing restrictions on the domain of admissible profiles. Kehne et al. (2025) shifted the focus on fairness to be more candidate-centric. Their desideratum of fairness states that candidates with similar sets of supporters should receive similar representation.
@@
Lastly, we mention the proportionality degree (PD) as an alternative for quantifying proportionality in multiwinner approval voting. This approach strictly enforces the quota while allowing greater flexibility in determining each group’s average satisfaction level. Skowron (2021) used this to measure the proportionality guarantees that different voting rules can achieve. He showed that the sequential variant of the proportional approval voting rule has higher PD than the sequential Phragmén rule. Janeczko and Faliszewski (2022) looked into computational aspects of the proportionality degree in committee elections. They show that deciding if a committee with a given PD exists is NP-hard and that verifying if a given committee provides a given PD is coNP-complete.
@@ 3 Preliminaries
We consider the standard approval-based multiwinner voting framework. For $x\in\operatorname{\mathbb{N}}$, let $x]=\{1,2,\dots,x\}$. Let $N=\{v_{1},\dots,v_{n}\}$ be the set of voters, $C=\{c_{1},\dots,c_{m}\}$ be the set of candidates, and $k\in\operatorname{\mathbb{N}}$ be the committee size with $k\leq m$. For every voter $v\in N$, we denote the set of candidates she approves of as $A_{v}\subseteq C$, and we write $N_{c}=\{v\in N:c\in A_{v}\}$ for the set of supporters of a candidate $c\in C$. A voting instance $I$ is defined by an approval profile $A=(A_{v})_{v\in N}$ and the committee size $k$, i.e., $I=(A,k)$. Finally, a committee $W$ is a subset of the candidates with cardinality $k$. To determine which committee to select, various multiwinner voting rules have been proposed. In this paper, we consider commonly used approval-based multiwinner voting rules, including the method of equal shares (MES), the sequential Phragmén (seq-Phragmén), the Chamberlin-Courant (CC), proportional approval voting (PAV), and the greedy justified candidate rule (GJCR). We refer the reader to the full version of the paper for the formal definitions of these rules.
@@ 4 Quota-Based Strengthening of Proportionality
As stated in the beginning, most proportionality axioms rely on the principle that a group is entitled to representation only if it satisfies some cohesiveness condition. Cohesiveness in its general form demands that a voter group satisfy the Hare quota (Aziz et al., 2017; Sánchez-Fernández et al., 2017).
@@
For $\ell\in[k]$, a group $S\subseteq N$ is $\ell$-cohesive if (1) $\lvert S\rvert\geq\ell\cdot\frac{n}{k}$ and (2) $\lvert\bigcap_{v\in S}A_{v}\rvert\geq\ell$.
In this work, we study a more general version of this definition, which incorporates a parameter $\alpha$ to make the quota requirement dynamic. The parameter $\alpha$ acts as a scaling factor that adjusts the group size required for representation. Smaller $\alpha$ values correspond to a more demanding standard of proportionality, as even smaller groups can claim representation.
@@
For $\ell\in[k]$ and $\alpha\in\operatorname{\mathbb{R}}_{\geq 0}$, a group $S\subseteq N$ is $(\alpha,\ell)$-cohesive if (1) $\lvert S\rvert\geq\alpha\cdot\ell\cdot\frac{n}{k}$ and (2) $\lvert\bigcap_{v\in S}A_{v}\rvert\geq\ell$.
For $S=\emptyset$, it holds that $\bigcap_{v\in S}A_{v}=C$. Changing the quota size has, up until now, mostly been done for $\alpha\geq 1$, which relaxes the cohesiveness definition (Dong et al., 2025a, b; Jiang et al., 2020; Do et al., 2022; Halpern et al., 2023).
With this generalized cohesiveness definition, we modify existing proportionality axioms:
@@
A committee $W$ satisfies $\alpha$-EJR+ if there does not exist a candidate $c\in C\setminus W$, a group of voters $S\subseteq N$, and $\ell\in[k]$ with $|S|\geq\alpha\cdot\ell\cdot\frac{n}{k}$ such that
@@
Clearly, for the case $\alpha=1$, the three proportionality axioms correspond to the original notions of JR, EJR, and EJR+. Therefore, several voting rules are known to guarantee satisfiability. More recently, it has been established that for any given instance, there exists an $\alpha<1$ such that $\alpha$-EJR+ can be satisfied.
@@
In this paper, we are interested in the minimal achievable quota for a given instance $I$. Therefore, we would like to find the infimum $\alpha$-value that still allows for an instance to satisfy $\alpha$-$\Phi$ where $\Phi\in\{\text{JR},\text{EJR},\text{EJR+}\}$. That being said, let
@@
for the given instance $I$. We refer to $\alpha_{\Phi}^{W}(I)$ as the $\alpha$-value of a committee $W$ with respect to the proportionality axiom $\Phi$. The minimum achievable $\alpha$-value for an instance $I$ is denoted by
$\alpha^{*}_{\Phi}(I)=\min_{W\subseteq C:\lvert W\rvert=k}\alpha^{W}_{\Phi}(I).$
We also refer to this as the optimal $\alpha$-value for an instance $I$ with respect to $\Phi$. If the instance $I$ is clear from the context, we omit it and just write $\alpha_{\Phi}^{*}$ or $\alpha_{\Phi}^{W}$.
[…]
If all voters with a non-empty ballot approve a candidate in the committee, we have $\alpha^{*}_{\text{JR}}=0$, and if $\lvert A_{v}\cap W\rvert\geq\min(\lvert A_{v}\rvert,k)$ for all $v\in N$, then $\alpha^{*}_{\text{EJR}}=0$.  [^3] For $\alpha>0$, every voter group $S$ with $\lvert S\rvert\geq\alpha\ell\frac{n}{k}$ has at least one ”satisfied” voter. Note that $\alpha\geq 0$ since the proportionality axioms are always violated when $\alpha=0$. For $\alpha>\frac{k}{k+1}$, all proportionality axioms can be satisfied, which was shown by (Casey and Elkind, 2025).